07-28-2014 04:54 PM
Have a pair pf HP Brocade 8/24 SAN switches (Brocade 5480) for the HP BladeSystem. For reasons I have yet to figure out one of the switches has decided to become unreachable via telnet or the web-based management URL.
At this time the management URL cannot be set via either EBIPA or manually via the OA CLI. In addition, the serial route to the OA is now disabled - so telnet is out as well.
All health indicators are good though I see that the Health LED has failed. HP's documentation says nothing about this particular error - or if it does I have yet to find any mention of it. With no information I as to what that partiular error means I can't say that the Health LED is dead since it is currently amber.
It is almost behaving as though it is somehow "lost" FOS though how I have no idea and even less if that's the case of how to restore/reload the OS given that I can't get in to the switch via either method.
Waiting on HP Support to get back to me but was wondering if anyone else has come across this set of problems and if so was there any solution?
Solved! Go to Solution.
08-06-2014 03:14 AM
After all the troubleshooting and efforts to get in to the switch it turned out that the management processor on the switch had failed.
I have a Care Pack on the switch so within 20 minutes of filing the case (at 7:00pm I might add) I got a call and was on the phone with an HP Tech who confirmed the troubleshooting steps and diagnosis and arranged to have a new switch shipped. Twenty minutes after hanging up the phone with the Tech I received a second call from HP's shipping center informing me that they were having the replacement couriered to me rather than the usual UPS delivery and would have the switch to me between 9 and 11am the following morning and telling me that I could have had it there earlier if I wished.
In the end all is well and as usual the service from HP was great.
08-06-2014 05:41 AM
Yes, tried both reseating the switch in the original interconnect bay (Bay 4) and also tried trading places with its partner switch in Bay 3 but that didn't resolve anything. I also restarted both OAs several times to see if that would do anything to reset the management processor but the issues remained.